. Construct a confidence interval of the population proportion at the given level of confidence.
x =300, n = 1200, 90% confidence.
Answer #1

Solution :

n = 1200

x = 300

\hat p = x / n =300 / 1200 = 0.250

1 -\hat p = 1 - 0.250 = 0.750

At 90% confidence level the z is ,

\alpha = 1 - 90% = 1 - 0.90 = 0.10

\alpha / 2 = 0.10 / 2 = 0.05

Z\alpha/2 = Z0.05 = 1.645

Margin of error = E = Z\alpha / 2 * \sqrt((\hat p * (1 - \hat p)) / n)

= 1.645 * (\sqrt((0.250 * 0.750) / 900)  

= 0.021

A 90 % confidence interval for population proportion p is ,

\hat p- E < P <\hat p + E

0.250 - 0.021 < p < 0.250 + 0.021

0.229 < p < 0.271